PLEASE HELP!!! At Backintime Cinema, tickets for adults cost $5.75 and tickets for children cost $3.75. How many of each kind of ticket were purchased if 16 tickets were bought for $68?
Answer by checkley71(8405)

5.75A+3.75C=68
A+C=16 OR A=16-C NOW REPLACE A WITH (16-C) IN THE FIRST EQUATION & SOLVE FOR C.
5.75(16-C)+3.75C=68
92-5.75C+3.75C=68
-2C=68-92
-2C=-24
C=-24/-2
C=12 CHILDREN TICKETS WERE SOLD.
16-12=4 ADULT TICKETS WERE SOLD.
PROOF:
5.75*4+3.75*12=68
23+45=68
68=68